# FAIRGAME: a Framework for AI Agents Bias Recognition using Game Theory

### v2

@@ 1 Introduction
AI agents powered by Large Language Models (LLM) are increasingly used in research [37], social [53] and industrial applications [45, 51], calling for the development of accurate prediction frameworks for their behaviors during interactions among themselves or with humans. Reliable predictions are essential for developing novel applications, promoting trustworthy AI systems, and mitigating undesirable outcomes and biases [10, 24]. Numerous approaches have been developed to improve the transparency and interpretability of individual AI agents [4, 19], as well as to identify their inconsistencies, biases and hallucinations [11, 35, 36]. However, less in known about cases where multiple interacting agents are involved [26], where emerging biases may skew strategic outputs in unpredictable manners. Studies emulating human behaviors [44] may produce spurious predictions. Also, in industry applications such as automated dispute resolution [9, 20], auction design or pricing mechanisms in finance and economics [7, 16, 17], or supply chain negotiations [2, 39, 47], hidden biases may result in unjust decisions, disproportionate favoring of certain groups, and distortion of fair competition.
@@
To address multi-agent interactions, and in addition to methods tailored to individual agents, game theory [43] formalizes interactions as games, so as to model, predict and optimize the strategic responses of rational agents [8, 20, 28]. Game theory has been employed to model and understand human choices in various contexts [50, 52], and AI-based players have been increasingly tested to reproduce classical game scenarios and provide additional complexity to them [15, 23, 56, 57], as well as to interact in game-like distributed technologies [30]. However, due to varying research protocols and discipline constraints, bridging the gap between theoretical game theory and empirical investigations of LLM agents in a reproducible, systematic and user-friendly setting is still a challenge [38].
@@
To facilitate seamless and reproducible integration of game-theoretic evaluations of AI interactions, we introduce FAIRGAME (Framework for AI Agents Bias Recognition using Game Theory), a versatile framework designed to simulate diverse scenarios, ranging from classical game theory models to realistic industrial use cases. In FAIRGAME, AI agents can be equipped with distinct features, such as strategic attitudes and personalities, linguistic variations, cultural orientations and more. The framework allows quantitative simulations of arbitrarily complex games in a systematic and reproducible manner, and to observe the emerging outcomes brought about by strategic interactions [29, 27], enabling direct comparison with game-theoretic predictions and supporting the inference of strategies through observations of multiple experiments [40]. Incorporating AI agents into controlled and predictable scenarios will also help identify and mitigate hidden biases related to language, cultural attributes, and more, which could result in suboptimal outcomes, unjust advantages, ethical dilemmas, or systemic inefficiencies [14, 22, 25].
@@ 2 Methodology: introducing FAIRGAME
FAIRGAME is a computational framework interfacing user-defined instructions to create the desired agents, eventually delivering standardized outputs for subsequent processing (Fig. 1). It allows us to test user-defined games, described textually via prompt injection and including any desired payoff matrix, as well as to define traits of the agents. The agents can be built from any LLM of choice, by calling the corresponding APIs (several of which are already provided in our package; however, we do not deal with payments of proprietary LLMs). FAIRGAME requires the following inputs. First, a Configuration File: A JSON file that defines the setup of both the agents and the game, in terms of parameters, payoff matrix entries, and additional information for the agents. For instance, agents can be associated with certain personalities, so as to increase the complexity of human behavior emulators [21, 31, 42] and predict the responses of personality-driven autonomous agents [34, 41]. Table 1 provides a detailed overview of the fields of the configuration file, along with their explanation. An example of configuration file is in Supplementary Section S1. And a Prompt Template: A text file that defines the instruction template, providing a literal description of the game. It includes the instructions for each agent, with each round’s parameters filled in from the configuration file, allowing customization of each agent. At runtime, this template is transformed into a prompt that includes details on available strategies, the corresponding payoffs based on decisions, and other configuration-specified factors – such as an agent’s personality and awareness of the opponent’s personality – as well as prior history information in case of repeated games. This file can be translated in any language of choice, allowing for multilingual tests. An example of an English prompt template for a Prisoner’s Dilemma game is provided in Supplementary Section S2. More than one prompt template can be associated to the same config file.
@@
opponentPersonalityProb List of Integers Probability that a certain agent has a certain personality, as referred to the other agents (see Supplementary Sec. A). If the probability is 0, this information is omitted from the prompt.
@@ 2.1 Creation and execution of games
Algorithm 1 processes the configuration file $CF$ and the set of prompt templates $PT$ as inputs, and outputs a list $G$ containing all instantiated games. The process involves validating inputs, extracting relevant game information, configuring agents, and finally creating the games.
@@
1:$CF$: Configuration file, $PT$: Prompt templates
@@
3:validate_config_file($CF$)
4:validate_templates($PT$, $CF$)
5:game_info, langs, llm, all_agent_perm $\leftarrow$ extract_info($CF$)
@@
7:    agents_config $\leftarrow$ compute_agents_combos($CF$, langs)
@@
9:    agents_config $\leftarrow$ get_agents_config($CF$, langs)
@@
11:$G$leftarrow ()
12:for $ac$ in agents_config do
13:    agents $\leftarrow$ create_agents($ac$, llm)
14:    game $\leftarrow$ create_games(game_info, agents, $PT$)
@@
[…]
@@
6:    prompt $\leftarrow$ create_prompt(template_lang, $g$.n_rounds, $g$.current_round, $g$.n_rounds_known, $g$.payoff_matrix, $g$.history()) )
@@ 3.1 LLMs and languages
We test AI agents from four widely used and publicly available large language models, as outlined in Table 2 together with their key details. We use default settings recommended by the providers.
Our study is conducted in five different written languages: English, French, Arabic, Vietnamese and Mandarin Chinese, to represent a diverse set of linguistic and cultural contexts, covering different scripts, grammatical structures, and global regions. This selection ensures a balanced and comprehensive analysis of language biases across widely spoken and culturally dominant languages. The template for each game is translated from English into each of the five target languages first by using an automated translator (see details in Supplementary Material Section S3), and then edited manually by a native speaker.
@@ 3.2 Games
We input their description in the template file, using standard game-theoretic narrative [43] (an example for the Prisoner Dilemma is in Supplementary Section S2). Each game is associated with a payoff matrix that represents the penalties incurred by various strategic choices, with player payoffs or gains expressed as negative values of these penalties. The matrices are in the form given by Table 3, and is parsed to the config file as described above.
@@
To explore strategic variations in adversarial interactions, and assess the sensitivity of each LLM to game parameters, we define multiple configurations of the Prisoner’s Dilemma game. Using an established scaling of dilemma strength [55], fixing other payoff matrix entries, the dilemma strength in the Prisoner’s Dilemma decrease with the difference between mutual reward and mutual punishment. For the conventional configuration with penalties $x_{1,1}=(6,6)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(2,2)$, this difference is $-2-(-6)=4$. The harsh configuration, with $x_{1,1}=(8,8)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(5,5)$ and the mild configuration, with $x_{1,1}=(8,8)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(2,2)$, have differences of $3$ and $6$, respectively.
For the Battle of the Sexes, we use a single configuration – the most commonly adopted in literature – with payoff matrix entries (referring to payoffs): $x_{1,1}=(10,7)$, $x_{1,2}=(0,0)$, $x_{2,1}=(0,0)$ and $x_{2,2}=(7,10)$.
@@ 3.3 Set up
The experimental configuration employed in this study are as follows. The name of each round depends on the game. The experiments consist of repeated games of 10 rounds each, without earlier stopping condition, for each LLM described in Tab. 2. We test both scenarios where agents are explicitly informed about the total number of rounds, and another where this information is withheld, as this might affect the outcome of the game theoretical predictions [6]. Tested languages are: [’en’, ’fr’, ’ar’, ’zh’, ’vn’]. We explicitly test the impact of agents’ personalities; here, we use ’cooperative’ and ’selfish’ (future works may even embed the OCEAN framework, or others [32]), to reflect general behavioral attributes commonly utilized and thoroughly documented in game theory literature. Conversely, agent identifiers were intentionally neutral (’agent1’ and ’agent2’) to eliminate additional variables that could introduce deviations from default behaviors, potentially compromising result interpretability (cf. Sec. 4.4). Personality traits are accurately translated into all evaluated languages, whereas agent identifiers remained untranslated, functioning purely as neutral placeholders. Importantly, agents are unaware of their opponent’s personality, a condition enforced through setting opponentPersonalityProb $=0$. All personality permutations were systematically generated, creating scenarios where both agents are cooperative, both selfish, or mixed configurations (one cooperative and one selfish). The payoffs are tailored for each game type, as described above. No early stopping condition was implemented: all 10 rounds are completed fully. Agents were not permitted to communicate during these experiments, leaving exploration of inter-agent interactions for future research.
The set of all configurations yields 18 distinct games per LLM. Each game is further tested 10 times to ensure statistical robustness. Considering 4 LLM, 5 languages, 10 rounds per game, and 2 decisions per round (one per agent), the experiment generated a total of 72,000 individual decisions.
@@ 4.1 Prisoner’s dilemma
Fig. 2 shows the bar plot (with $95\%$ CI) summarizing the test results (in terms of final penalties received by the agents) of the Prisoner’s Dilemma games for all three versions (a breakdown for each version, which maintains alike patterns, is reported in Supplementary Section S4) and under two conditions: when agents are unaware of their opponent’s personality and when they are informed. The results are shown across all considered LLMs and languages examined in this study, and for all personality combinations.
In most cases, the preferred end result favors agents defecting, as suggested by game theory. This aligns with the Nash equilibrium of the Prisoner’s Dilemma, where mutual defection is the dominant strategy. Nevertheless, there are inconsistencies across languages and combinations of personalities, which suggests that the agents’ behavior is influenced by factors beyond the payoff matrix, such as languages and inherent biases present in LLM training data. For instance, penalties are generally lower in English, particularly for GPT-4o and Claude 3.5 Sonnet, when the number of rounds is unknown, indicating more consistent cooperative behavior in their primary training language. Broader variability in languages like French or Arabic with high penalties suggests challenges in interpreting the game dynamics due to linguistic or cultural differences, while penalties remain high in Chinese and Vietnamese, particularly for Mistral Large. In mixed personality settings (CS), selfish agents exploit cooperative ones, leading to higher penalties for the latter, consistent with game-theoretic predictions. For selfish pairings (SS), penalties are high but exhibit low variability, as mutual defection is the rational choice. When the number of rounds is unknown, penalties are higher across all settings, reflecting uncertainty that discourages cooperation. Conversely, when rounds are known, penalties decrease, particularly in CC and CS settings, as agents can plan strategies with the endgame in mind. Claude 3.5 Sonnet and GPT-4o show significant reductions in penalties when rounds are known, especially in English, demonstrating their ability to adapt to game structure. In contrast, Mistral Large shows less sensitivity to this information. Overall, knowing the number of rounds promotes cooperation in CC and CS settings, while SS settings remain unaffected, as defection remains the dominant strategy.
@@
Figure 3: Evolution of normalized penalties (averaged over repeated experiments) over repeated rounds, for each LLM and variations of the Prisoner’ dilemma.
@@
We also study the evolution of strategies over the rounds. In Fig. 3, the average outcomes obtained by the LLM at each round are normalized (to facilitate direct comparison across configurations with different payoff matrices) to a scale ranging from -1 to 1, corresponding respectively to the minimum and maximum achievable penalties in each game. We show the evolution for each version of the game (cf. Sec. 3.2). Tuning the payoffs changes the outcomes of the games: all LLMs exhibit, on average, higher penalties under the harsh scenario and lower penalties under the mild scenario compared to the conventional scenario, consistent with the game-theoretic principles of repeated games [55]. In the harsh scenario, higher penalties discourage cooperation, leading to more defection. Conversely, the mild scenario incentivizes cooperation, resulting in lower penalties over time, particularly for Claude 3.5 Sonnet and Llama 3 405B. The conventional scenario balances these effects, with penalties stabilizing at intermediate levels. We also observe the variability between harsh and mild scenarios, indicative of each model’s sensitivity to payoff conditions when making decisions. Claude 3.5 Sonnet and Llama 3 405B demonstrate lower sensitivity, reflected by a narrower spread, while Mistral and GPT-4o show higher sensitivity to the parameters. Llama 3 405B and GPT-4o exhibit less differentiation between the conventional and mild scenarios, whereas Mistral displays smaller differences between harsh and conventional conditions. This behavior suggests distinct decision-making strategies influenced by payoff variations. Finally, we observe a general downward trend in penalties over time for Claude 3.5 and Llama 3 405B, indicating progressively increasing mutual cooperation among agents, consistent with reciprocal strategies in repeated games, where agents reciprocate cooperation to maximize long-term payoffs [6, 55]. Conversely, Mistral Large shows stable cooperation levels under conventional and harsh scenarios but a marked increase in cooperation within the mild scenario. GPT-4o exhibits divergent patterns, with decreasing penalties (increasing cooperation) in the harsh scenario and increasing penalties (decreasing cooperation) in both mild and conventional scenarios. This divergent behavior reflects context-dependent strategic adaptation, potentially due to its higher variability in interpreting payoff structures. These results highlight the interplay between payoff matrices and strategic behavior in repeated games.

### v3

@@ 1 Introduction
AI agents powered by Large Language Models (LLM) are increasingly used in research [36], social [52] and industrial applications [44, 50], calling for the development of accurate prediction frameworks for their behaviors during interactions among themselves or with humans. Reliable predictions are essential for developing novel applications, promoting trustworthy AI systems, and mitigating undesirable outcomes and biases [11, 23]. Numerous approaches have been developed to improve the transparency and interpretability of individual AI agents [5, 19], as well as to identify their inconsistencies, biases and hallucinations [12, 33, 34]. However, less is known about cases where multiple interacting agents are involved [25], where emerging biases may skew strategic outputs in unpredictable manners. Studies emulating human behaviors [43] may produce spurious predictions. Also, in industry applications such as automated dispute resolution [10, 20], auction design or pricing mechanisms in finance and economics [8, 16, 17], or supply chain negotiations [3, 38, 46], hidden biases may result in unjust decisions, disproportionate favoring of certain groups, and distortion of fair competition.
@@
To address multi-agent interactions, and in addition to methods tailored to individual agents, game theory [42] formalizes interactions as games, so as to model, predict and optimize the strategic responses of rational agents [9, 20, 26]. Game theory has been employed to model and understand human choices in various contexts [49, 51], and AI-based players have been increasingly tested to reproduce classical game scenarios and provide additional complexity to them [22, 55, 56], as well as to interact in game-like distributed technologies [28]. However, due to varying research protocols and discipline constraints, bridging the gap between theoretical game theory and empirical investigations of LLM agents in a reproducible, systematic and user-friendly setting is still a challenge [37].
To facilitate seamless and reproducible integration of game-theoretic evaluations of AI interactions, we introduce FAIRGAME (Framework for AI Agents Bias Recognition using Game Theory), a versatile framework designed to simulate diverse scenarios, ranging from classical game theory models to realistic industrial use cases. FAIRGAME is an open-source project [1].
@@
In FAIRGAME, AI agents can be equipped with distinct features, such as strategic attitudes and personalities, linguistic variations, cultural orientations and more. The framework allows quantitative simulations of arbitrarily complex games in a systematic and reproducible manner, and to observe the emerging outcomes brought about by strategic interactions [27, 13], enabling direct comparison with game-theoretic predictions and supporting the inference of strategies through observations of multiple experiments [39]. Incorporating AI agents into controlled and predictable scenarios will also help identify and mitigate hidden biases related to language, cultural attributes, and more, which could result in suboptimal outcomes, unjust advantages, ethical dilemmas, or systemic inefficiencies [15, 21, 24].
@@ 2 Methodology: introducing FAIRGAME
FAIRGAME is a computational framework interfacing user-defined instructions to create the desired agents, eventually delivering standardized outputs for subsequent processing (Fig. 1). It allows to test user-defined games, described textually via prompt injection and including any desired payoff matrix, as well as to define traits of the agents. The agents can be built from any LLM of choice, by calling the corresponding APIs (several of which are already provided in our package; however, any usage fees are covered by the end user). FAIRGAME requires the following inputs. First, a Configuration File: a JSON file that defines the setup of both the agents and the game, in terms of parameters, payoff matrix entries, and additional information for the agents. For instance, agents can be associated with certain personalities, so as to increase the complexity of human behavior emulators [29, 41] and predict the responses of personality-driven autonomous agents [32, 40]. Table 1 provides a detailed overview of the fields of the configuration file, along with their explanation. An example of configuration file is in Supplementary Material Section S1 [2]. APrompt Template: a text file that defines the instruction template, providing a literal description of the game. It includes the instructions for each agent, with each round’s parameters filled in from the configuration file, allowing customization of each agent. At runtime, this template is transformed into a prompt that includes details on available strategies, the corresponding payoffs based on decisions, and other configuration-specified factors – such as an agent’s personality and awareness of the opponent’s personality – as well as prior history information in case of repeated games. This file can be translated in any language of choice, allowing for multilingual tests. An example of an English prompt template for a Prisoner’s Dilemma game is provided in Supplementary Material Section S2 [2]. More than one prompt template can be associated to the same config file.
@@
opponentPersonalityProb List of Integers Probability that a certain agent has a certain personality, as referred to the other agents. If the probability is 0, this information is omitted from the prompt.
@@ 2.1 Creation and execution of games
Algorithm 1 processes the configuration file CF and the set of prompt templates $PT$ as inputs, and outputs a list $G$ containing all instantiated games. The process involves validating inputs, extracting relevant game information, configuring agents, and creating the games.
@@
1:CFCF: Configuration file, PTPT: Prompt templates
@@
3:validate_config_file(CFCF)
4:validate_templates(PTPT, CFCF)
5:game_info, langs, llm, all_agent_perm $\leftarrow$ extract_info(CFCF)
@@
7:  agents_config $\leftarrow$ compute_agents_combos(CFCF, langs)
@@
9:  agents_config $\leftarrow$ get_agents_config(CFCF, langs)
@@
11:$G$ $\leftarrow$ ()
12:for acac in agents_config do
13:  agents $\leftarrow$ create_agents(acac, llm)
14:  game $\leftarrow$ create_games(game_info, agents, PTPT)
@@
Once instantiated, games are executed as per Algorithm 2. An empty list $O$ is initialized (step 1), which will be populated with the outcomes of all games after execution. The list $G$ is then given as input to the algorithm. Each game $g$ in $G$ is processed independently (steps 2–9). If the user requests multiple rounds in an evolutionary game theory setting, a while loop governs their execution. The loop continues as long as two conditions are satisfied: the current round count does not exceed the maximum number of rounds $g.\textit{n\_rounds}$, and the game-specific stopping condition $g.\textit{stop\_cond}$ is not met (step 4). The stopping condition allows for the game to terminate early based on predefined criteria, ensuring flexibility in simulation. Within each iteration of the while loop, the function $g.\textit{run\_round()}$ is called to execute the logic for the current round (step 5). This function, which is described in detail in Algorithm 3, simulates the actions of the agents involved in the game. Note that communication between agents is supported by FAIRGAME; however, we do not use it in the next use cases, hence the algorithm described here represents a simplified version that does not take inter-agent communication into account. We have dedicated a separate study to analyzing the effect of communication by comparing outcomes with and without it [14].
Once the game terminates -— either because the maximum number of rounds has been reached or the stopping condition is satisfied -— the algorithm retrieves the game’s history, which is appended to the output file $O$ (step 8).
@@
6:  prompt $\leftarrow$ create_prompt(template_lang, $g$.n_rounds, $g$.current_round, $g$.n_rounds_known, $g$.payoff_matrix, $g$.history())
@@ 3.1 LLMs and languages
We evaluate AI agents using four widely-used and publicly available LLMs, described along with their key details in Table 2. All models were tested using the default settings recommended by their respective providers. For all LLMs, we used the latest available versions at the time of the experiments, conducted from 10 to 15 February 2025.
Our study is conducted in five different written languages: English, French, Arabic, Vietnamese and Mandarin Chinese, to represent a diverse set of linguistic and cultural contexts, covering different scripts, grammatical structures, and global regions. This selection ensures a balanced and comprehensive analysis of language biases across widely spoken and culturally dominant languages. The template for each game is translated from English into each of the five target languages first by using an automated translator (see details in Supplementary Material Section S3 [2]), and then edited manually by a native speaker. The personality traits were also revised by native speakers.
@@ 3.2 Games
We input their description in the template file, using standard game-theoretic narrative [42] (an example for the Prisoner Dilemma is in Supplementary Material Section S2 [2]). Each game is associated with a payoff matrix that represents the penalties or rewards incurred by various strategic choices, with player payoffs or gains expressed as negative values of these penalties. The matrices are in the form given by Table 3, and is parsed to the config file as described above.
@@
To explore strategic variations in adversarial interactions, and assess the sensitivity of each LLM to game parameters, we define multiple configurations of the Prisoner’s Dilemma game. Using an established scaling of dilemma strength [54], fixing other payoff matrix entries, the dilemma strength in the Prisoner’s Dilemma decrease with the difference between mutual reward and mutual punishment. For the conventional configuration with penalties $x_{1,1}=(6,6)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(2,2)$, this difference is 2(6)=4-2-(-6)=4. The harsh configuration, with $x_{1,1}=(8,8)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(5,5)$ and the mild configuration, with $x_{1,1}=(8,8)$, $x_{1,2}=(0,10)$, $x_{2,1}=(10,0)$ and $x_{2,2}=(2,2)$, have differences of $3$ and $6$, respectively.
The Battle of the Sexes uses a single (the most commonly adopted in literature) configuration, with matrix entries (referring to payoffs): $x_{1,1}=(10,7)$, $x_{1,2}=(0,0)$, $x_{2,1}=(0,0)$ and $x_{2,2}=(7,10)$.
@@ 3.3 Set up
The experimental configuration employed in this study is as follows. The name of each round depends on the game. The experiments consist of repeated games of 10 rounds each, without earlier stopping condition, for each LLM described in Tab. 2. We test both scenarios where agents are explicitly informed about the total number of rounds, and another where this information is withheld, as this might affect the outcome of the game theoretical predictions [7]. Tested languages are: [’en’, ’fr’, ’ar’, ’zh’, ’vn’]. We explicitly test the impact of agents’ personalities; here, we use ’cooperative’ and ’selfish’ (future works may even embed the OCEAN framework, or others [30]), to reflect general behavioral attributes commonly utilized and thoroughly documented in game theory literature. Conversely, agent identifiers were intentionally neutral (’agent1’ and ’agent2’) to eliminate additional variables that could introduce deviations from default behaviors, potentially compromising result interpretability (cf. Sec. 4.4). Personality traits are accurately translated into all evaluated languages, whereas agent identifiers remained untranslated, functioning purely as neutral placeholders. Importantly, agents are unaware of their opponent’s personality, a condition enforced through setting opponentPersonalityProb =0=0. All personality permutations were systematically generated, creating scenarios where both agents are cooperative, both selfish, or mixed configurations (one cooperative and one selfish). The payoffs are tailored for each game type, as described above. No early stopping condition was implemented: all 10 rounds are completed fully. Agents were not permitted to communicate during these experiments, leaving exploration of inter-agent interactions for future research.
The set of all configurations yields 18 distinct games per LLM. Each game is further tested 10 times to ensure statistical reliability. Considering 4 LLM, 5 languages, 10 rounds per game, and 2 decisions per round (one per agent), the experiment generated a total of 72,000 individual decisions.
@@ 4.1 Prisoner’s dilemma
Fig. 2 shows the bar plot (with $95\%$ Confidence Interval) summarizing the test results (in terms of final penalties received by the agents) of the Prisoner’s Dilemma games for all three versions (a breakdown for each version, which maintains alike patterns, is reported in Supplementary Material Section S4), and under two conditions: when agents are unaware of their opponent’s personality and when they are informed. The results are shown across all considered LLMs and languages examined in this study, and for all personality combinations.
In most cases, the preferred end result favors agents defecting, as suggested by game theory. This aligns with the Nash equilibrium of the Prisoner’s Dilemma, where mutual defection is the dominant strategy. Nevertheless, there are inconsistencies across languages and combinations of personalities, which suggests that the agents’ behavior is influenced by factors beyond the payoff matrix, such as languages and inherent biases present in LLM training data. For instance, penalties are generally lower in English, particularly for GPT-4o and Claude 3.5 Sonnet, when the number of rounds is unknown, indicating more consistent cooperative behavior in their primary training language. Broader variability in languages like French or Arabic with high penalties suggests challenges in interpreting the game dynamics due to linguistic or cultural differences, while penalties remain high in Mandarin Chinese and Vietnamese, particularly for Mistral Large. In mixed personality settings (CS), selfish agents exploit cooperative ones, leading to higher penalties for the latter, consistent with game-theoretic predictions. For selfish pairings (SS), penalties are high but exhibit low variability, as mutual defection is the rational choice. When the number of rounds is unknown, penalties are higher across all settings, reflecting uncertainty that discourages cooperation. Conversely, when rounds are known, penalties decrease, particularly in CC and CS settings, as agents can plan strategies with the endgame in mind. Claude 3.5 Sonnet and GPT-4o show significant reductions in penalties when rounds are known, especially in English, demonstrating their ability to adapt to game structure. In contrast, Mistral Large shows less sensitivity to this information. Overall, knowing the number of rounds promotes cooperation in CC and CS settings, while SS settings remain unaffected, as defection remains the dominant strategy.
@@
Figure 3: Average trajectory of strategy choices across repeated rounds in all Prisoner’s Dilemma experiments, presented for each LLM and game variant. A value of 1 denotes selection of Option A, which corresponds to defection in this game, while -1 represents Option B (cooperation).
@@
We also study the evolution of strategies over the rounds. In Fig. 3, we show the evolution for each version of the game (cf. Sec. 3.2). The figure shows that tuning the payoffs changes the strategies adopted during repeated games: all LLMs exhibit, on average, more selfishness under the harsh scenario and more cooperation under the mild scenario compared to the conventional scenario, consistent with the game-theoretic principles of repeated games [54]. In the harsh scenario, higher penalties discourage cooperation, leading to more defection. Conversely, the mild scenario incentivizes cooperation particularly for Claude 3.5 Sonnet and Llama 3.1 405B. The conventional scenario balances these effects, with strategies stabilizing at intermediate levels. We also observe the variability between harsh and mild scenarios, indicative of each model’s sensitivity to payoff conditions when making decisions. Claude 3.5 Sonnet and Llama 3.1 405B demonstrate lower sensitivity, reflected by a narrower spread, while Mistral and GPT-4o show higher sensitivity to the parameters. Llama 3.1 405B and GPT-4o exhibit less differentiation between the conventional and mild scenarios, whereas Mistral displays smaller differences between harsh and conventional conditions. This behavior suggests distinct decision-making strategies influenced by payoff variations. Finally, we observe a general downward trend in selfishness over time for Claude 3.5 and Llama 3.1 405B, indicating progressively increasing mutual cooperation among agents, consistent with reciprocal strategies in repeated games, where agents reciprocate cooperation to maximize long-term payoffs [7, 54]. Conversely, Mistral Large shows stable cooperation levels under conventional and harsh scenarios but a marked increase in cooperation within the mild scenario. GPT-4o exhibits divergent patterns, with increasing cooperation in the harsh scenario and increasing selfishness in both mild and conventional scenarios. This divergent behavior reflects context-dependent strategic adaptation, potentially due to its higher variability in interpreting payoff structures. These results highlight the interplay between payoff matrices and strategic behavior in repeated games.